alexusH:

Mary walked north from her home to Sheila's home, which is 4.0 kilometers away. Then she turned right and walked another 3.0 kilometers to the supermarket, which is 5.0 kilometers from her own home. She walked the total distance in 1.5 hours. What were her average speed and average velocity?

justus:

The average speed is total distance travelled divided by the time: s = (d1 + d2)/t = 4+3/1.5 = 7/1.5 = The average velocity (v) is total displacement (d₁) from the point at the start divided by time. v = d₁/t = 5/1.5 =